# ArticulatePro: A Comparative Study on a Proactive and Non-Proactive Assistant in a Climate Data Exploration Task

### v1

@@ 3.1.1. Investigating Proactive Opportunities
To explore the various opportunities for a digital assistant to be proactive, we performed an analysis on our prior work (Tabalba et al., 2023b). In our prior work, we conducted a study to investigate the benefits of an always-listening digital assistant. The study involved dyads of participants interacting with a digital assistant for data exploration. This study allowed us to observe different opportunities a digital assistant could be proactive during a data exploration task. This was a necessary step for our research because there has been no prior work in the literature that focused on investigating proactive opportunities in a data exploration task. For our analysis, one researcher reviewed the study’s utterances and developed classifications for situations where a digital assistant has the potential to be proactive. Here are the classifications that we identified:
@@
Users tend to verbally state their findings or discoveries. For example, a user might examine a chart and say, ”So as COVID rates increase, so does poverty,” or ”Social vulnerability seems to have a lot in common with the map of poverty.” This may be an opportunity for the assistant to confirm or deny the user’s discovery.
@@
Users might express disagreement with each other, saying things like, ”I don’t think that is true,” or ”The chart doesn’t necessarily show that.” This may be an opportunity where the assistant provides data to settle the disagreement between users.
@@
Users might also state their preference for the type of chart they would like to see, such as, ”I think a good strategy is to bring up a lot of maps.” For this opportunity, the assistant could remember to stick to chart types that the user prefers.
@@
Users may criticize the charts that were given to them. These may be moments where the users state ”What does that say? I’m blind, I can’t see the text”. For this opportunity, the assistant could increase the font text size in the current chart and future charts.
@@
Users may preface the chart that they are going to create next by saying ”It seems like the Southeast and Small City would be good places to start. Let’s start by prioritizing resources based on these variables.” This could be an opportunity where the digital assistant creates charts on the areas the user is interested in before they ask for it.
@@
Users might be confused on the purpose of their task by saying ”When we talk about prioritizing resources, what are resources? Do doctors count as a resource?”. This could be a moment where the assistant could offer its own input.
These classifications we observed were also consistent with the envisioned proactive assistant storyboards that was brainstormed by Zargam et al and Reicherts et al. (Reicherts et al., 2021; Zargham et al., 2022). However, their classifications were based on the context of home assistants. Our classifications were based on the context of data exploration.
Out of the opportunistic proactive classifications that we identified, we wanted to focus our attention towards the most prevalent classification. Therefore we choose when users make a finding or discovery, which was the most prevalent classification that we found. By selecting this classification, we maximized our chances of observing user reactions to a proactive assistant, while limiting the complexity of our study by focusing on a single aspect of proactivity.
@@ 3.1.2. Pragmatic Approach for Proactive Opportunities
For our approach, we applied Pragmatic theories to guide the development of our proactive assistant. According to Grice’s principles of Cooperative Conversation (Grice, 1975), communication is most effective when it adheres to the principle of relevance (Wilson and Sperber, 1986). Based on this, we incorporated a ”history” component into our assistant, allowing it to track the conversation, the user’s interactions, and the charts it has already generated. This enables the system to produce relevant charts based on the ongoing dialogue and the user’s preferences while minimizing redundancy. For more details on our implementation oh history, see Section 3.2.3.
@@ 6.2. Enhanced Reliability of Findings
The proactive digital assistant, Arti, excelled at presenting the same data in different chart formats, which often helped users confirm or refute their hypotheses. For instance, in session 12, participants were analyzing a box plot showing rainfall for Kauai. Initially, they quickly concluded that Kauai received no rainfall. However, Arti proactively generated a scatter plot showing rainfall and air temperature. Initially confused by data indicating instances of rainfall, the participants re-evaluated their conclusion and realized they had misinterpreted the original chart. This example demonstrates how a proactive system can aid users in validating or challenging their hypotheses, aligning with Sperber and Wilson’s notion of ”stimulating the cognitive environment” (Wilson and Sperber, 1986). In Pragmatics, stimulating the cognitive environment refers to whether an utterance contributes meaningfully to the conversation. In contrast, when an utterance lacks usefulness, it may fail to spark interesting or productive thoughts.
This suggests that proactive systems have the potential to provide reassurance in the decisions and findings users make during data exploration. Similarly, Reicherts et al. developed ”ProberBot,” a chatbot designed to prompt users during investment decisions by engaging them in dialogue to justify their reasons for holding, buying, or selling stocks (Reicherts et al., 2022). The study found that ProberBot encouraged users to take more time to re-evaluate their decisions, which led to more logical choices and reduced cognitive and emotional biases.
However, Reicherts et al. also observed that participants felt the bot’s dialogue could bias their decisions. In our case, while Arti can challenge users’ hypotheses, we recognize the importance of ensuring that the assistant doesn’t influence users toward incorrect beliefs. Like Reicherts et al., we emphasize that AI assistants must remain neutral in decision-making tasks and avoid steering users’ thought processes.
@@ 6.3. Greater Usage of Different Chart Types
Arti broadened users’ exploration of chart types. A common approach among users was to rely on a single chart type, with line charts being the most prevalent. In session 2, participants mainly used line charts with Marti in the first session, briefly switching to histograms near the end. However, when they interacted with Arti in the second session, Arti quickly generated a box plot—something the users had not explored with Marti. They continued using box plots throughout the session. A similar effect was observed in session 6, where participants only used line charts with Marti but explored scatter plots, box plots, histograms, and line charts with Arti, greatly expanding their chart usage. This variety helped participants draw more insights from the data.
@@
In a previous study, Pins et al. examined how people interact with home voice assistants (Pins et al., 2020). They found that systems like Google Assistant and Amazon Alexa are often underutilized, with users sticking to simple commands, especially after failed interactions. One problem is that the assistants’ capabilities are either listed in a user manual or voiced upon request, making it difficult for users to remember them. In contrast, a visual display can remind users of a system’s capabilities. Our study shows the potential of proactive systems to suggest new features, helping users fully utilize the system’s capabilities.
6.4. User’s Acceptance of a Proactive Assistant
@@
We found that users did not prefer to start with the proactive digital assistant. As Figure 8 illustrates, none of the participants who began with the proactive assistant preferred to continue working with it. According to feedback from the post-interview, starting with Arti may have been too overwhelming for participants as they began the task. This is understandable given the multiple factors they had to manage: learning how to communicate with the digital assistant to create visualizations, coordinating with their partner on how to approach climate data, and doing so in an unfamiliar environment. In contrast, participants who started with Marti had a chance to familiarize themselves with the system and dataset before transitioning to the proactive assistant, making the shift to Arti easier for them.
@@
For those developing a proactive digital assistant for their own use cases, we advise caution in not overwhelming users. One participant noted that the proactive digital assistant was “too fast,” often creating useful charts before they even asked for them. While our findings suggest that starting with a proactive digital assistant might help users learn the system more quickly, it’s clear that users do not prefer to start with it. One potential avenue for future work would be investigate ways the assistant could adjust to the user’s level of proactivity. Prior studies have investigated ”context-aware” systems, where the computer has the ability to measure the user’s surroundings like computer vision, sensors, temperature, etc (Schmidt et al., 1999; Carvalho et al., 2017; Ferrero et al., 2019). For instance, one can imagine how using a camera that observes the user’s face, eyes, mouse movements, or body language to detect if the user has been able to absorb the information, before providing more.
@@ 7. Conclusion
In this study, we developed and evaluated a proactive data exploration assistant designed to generate relevant content and support users during data analysis tasks. Our approach was grounded in Pragmatics, which posits that effective utterances are those that relevant to the conversation. Our findings support this theory, as users made more task-related utterances and discoveries with the proactive assistant compared to the non-proactive one. Additionally, our results demonstrate how a proactive assistant can improve system learnability and effectiveness, improve the reliability of users’ findings, and increase engagement with the system’s charting capabilities.
While the results indicated that users may have the potential to be more efficient with the proactive assistant, as it helped them initiate analysis more quickly through automatic chart generation, this needs further investigation. Specifically, a between-subject design would be more suitable for thoroughly assessing the potential efficiency benefits.
Interestingly, while users were more efficient with the proactive assistant, many expressed a preference for less proactivity. This suggests a delicate balance between offering assistance and allowing users to maintain control over their workflow. When the proactive assistant was too intrusive, users tended to ignore its input in favor of their own strategies. However, dialing back proactivity may risk limiting the assistant’s ability to prompt new discoveries and guide users toward more diverse chart types.
Our findings suggest that the proactive assistant positively impacted both task efficiency and the depth of user exploration. As technology continues to advance, there is an opportunity to refine refine natural language systems to more closely emulate human communication, as described by linguistic theories like Pragmatics. This could enable more intuitive and seamless interactions between users and digital systems.

### v2

@@ 3.1.1. Investigating Proactive Opportunities
To explore the various opportunities for a digital assistant to be proactive, we conducted an analysis based on our prior work (Tabalba et al., 2023b). In that study, we investigated the benefits of an always-listening digital assistant by having dyads of participants interact with it during data exploration tasks. This study allowed us to observe potential moments where a digital assistant could be proactive. This step was essential for our research, as no prior work in the literature has focused on investigating proactive opportunities in data exploration tasks. For our analysis, one researcher reviewed the study’s utterances and developed classifications for situations where a digital assistant could potentially be proactive. The following classifications were identified:
@@
Users often verbally state their findings or discoveries. For example, a user might examine a chart and say, ”So as COVID rates increase, so does poverty,” or ”Social vulnerability seems to overlap with the map of poverty.” This presents an opportunity for the assistant to confirm or challenge the user’s discovery.
@@
Users might express disagreement with one another, saying things like, ”I don’t think that’s true,” or ”The chart doesn’t necessarily show that.” In such cases, the assistant could provide relevant data to resolve the disagreement.
@@
Users may state their preference for the type of chart they want to see, such as, ”I think a good strategy is to use a lot of maps.” Here, the assistant could remember the user’s preference and prioritize similar chart types in the future.
@@
Users may criticize the charts provided to them, saying things like, ”What does that say? I’m blind, I can’t read the text.” In such cases, the assistant could adjust the chart by increasing font size or making other enhancements.
@@
Users may indicate what they plan to explore next, saying something like, ”It seems like the Southeast and small cities would be good places to start. Let’s prioritize resources based on these variables.” This is an opportunity for the assistant to generate relevant charts based on the user’s interests before they explicitly ask for it.
@@
Users might express confusion about the task at hand, saying something like, ”When we talk about prioritizing resources, what are resources? Do doctors count as a resource?” In such moments, the assistant could step in to clarify or offer guidance.
These classifications are consistent with the proactive assistant storyboards envisioned by Zargham et al. and Reicherts et al. (Reicherts et al., 2021; Zargham et al., 2022). However, while their classifications were based on home assistants, ours are focused on the context of data exploration.
Out of the identified proactive classifications, we decided to focus on the most prevalent one: when users make a finding or discovery. This classification was the most common in our study and, by selecting it, we maximized the chances of observing user reactions to a proactive assistant while limiting the complexity of our study to a single aspect of proactivity.
@@ 3.1.2. Pragmatic Approach for Proactive Opportunities
To guide the development of our proactive assistant, we applied Pragmatic theories. According to Grice’s principles of Cooperative Conversation (Grice, 1975), communication is most effective when it adheres to the principle of relevance (Wilson and Sperber, 1986). Based on this, we incorporated a ”history” component into our assistant, enabling it to track the conversation, the user’s interactions, and the charts it has already generated. This allows the system to produce relevant charts based on the ongoing dialogue and user preferences while minimizing redundancy. For more details on our implementation of the history component, see Section 3.2.3.
@@ 6.2. Enhanced Reliability of Findings
The proactive digital assistant, Arti, excelled at presenting the same data in different chart formats, which often helped users confirm or refute their hypotheses. For example, in session 12, participants were analyzing a box plot showing rainfall for Kauai. Initially, they quickly concluded that Kauai received no rainfall. However, Arti proactively generated a scatter plot showing rainfall and air temperature. Initially confused by data indicating instances of rainfall, the participants re-evaluated their conclusion and realized they had misinterpreted the original chart. This example demonstrates how a proactive system can help users validate or challenge their hypotheses, aligning with Sperber and Wilson’s notion of ”stimulating the cognitive environment” (Wilson and Sperber, 1986). In Pragmatics, stimulating the cognitive environment refers to how an utterance contributes meaningfully to the conversation. In contrast, when an utterance lacks usefulness, it may fail to spark interesting or productive thoughts.
This suggests that proactive systems have the potential to reassure users in the decisions and findings they make during data exploration. Similarly, Reicherts et al. developed ”ProberBot,” a chatbot designed to prompt users during investment decisions by engaging them in dialogue to justify their reasons for holding, buying, or selling stocks (Reicherts et al., 2022). The study found that ProberBot encouraged users to take more time to re-evaluate their decisions, leading to more logical choices and reduced cognitive and emotional biases.
However, Reicherts et al. also observed that participants felt the bot’s dialogue could bias their decisions. In our case, while Arti can challenge users’ hypotheses, we recognize the importance of ensuring that the assistant does not influence users toward incorrect beliefs. Like Reicherts et al., we emphasize that AI assistants must remain neutral in decision-making tasks and avoid steering users’ thought processes.
@@ 6.3. Greater Usage of Different Chart Types
Arti broadened users’ exploration of chart types. A common approach among users was to rely on a single chart type, with line charts being the most prevalent. In session 2, participants primarily used line charts with Marti in the first session, briefly switching to histograms near the end. However, when they interacted with Arti in the second session, Arti quickly generated a box plot—something the users had not explored with Marti. They continued using box plots throughout the session. A similar effect was observed in session 6, where participants initially used only line charts with Marti but explored scatter plots, box plots, histograms, and line charts with Arti, greatly expanding their chart usage. This variety helped participants draw more insights from the data.
@@
In a previous study, Pins et al. examined how people interact with home voice assistants (Pins et al., 2020). They found that systems like Google Assistant and Amazon Alexa are often underutilized, with users sticking to simple commands, especially after failed interactions. One issue is that these assistants’ capabilities are either listed in a user manual or voiced upon request, making it difficult for users to remember them. In contrast, a visual display can remind users of a system’s capabilities. Our study shows the potential of proactive systems to suggest new features, helping users fully utilize the system’s capabilities.
6.4. Users’ Acceptance of a Proactive Assistant
@@
We found that users did not prefer to start with the proactive digital assistant. As illustrated in Figure 8, none of the participants who began with the proactive assistant preferred to continue working with it. According to feedback from post-interviews, starting with Arti may have been overwhelming for participants as they began the task. This is understandable given the multiple factors they had to manage: learning how to communicate with the digital assistant to create visualizations, coordinating with their partner on how to approach the climate data, and doing so in an unfamiliar environment. In contrast, participants who started with Marti had a chance to familiarize themselves with the system and dataset before transitioning to the proactive assistant, making the shift to Arti easier.
@@
For those developing a proactive digital assistant for their own use cases, we advise caution in not overwhelming users. One participant noted that the proactive digital assistant was “too fast,” often creating useful charts before they even asked for them. While our findings suggest that starting with a proactive digital assistant might help users learn the system more quickly, it’s clear that users do not prefer to start with it. Future work could investigate ways for the assistant to adjust its level of proactivity based on the user’s needs. Prior studies have explored ”context-aware” systems, where computers measure the user’s surroundings using technologies like computer vision, sensors, or temperature monitoring (Schmidt et al., 1999; Carvalho et al., 2017; Ferrero et al., 2019). For instance, one could imagine using a camera to observe the user’s face, eyes, mouse movements, or body language to detect whether the user has absorbed the information before providing more.
@@ 7. Conclusion
In this study, we developed and evaluated a proactive data exploration assistant designed to generate relevant content and support users during data analysis tasks. Our approach was motivated by Pragmatics, which posits that effective utterances are those that are relevant to the conversation. Our findings support this theory, as users made more task-related utterances and discoveries with the proactive assistant compared to the non-proactive one. Additionally, our results demonstrate how a proactive assistant can improve system learnability, enhance the reliability of users’ findings, and increase engagement with the system’s charting capabilities. While the results indicated that users could potentially be more efficient with the proactive assistant, as it helped them initiate analysis more quickly through automatic chart generation, further investigation is needed. Specifically, a between-subject design would be more suitable for thoroughly assessing the potential efficiency benefits. Interestingly, while users were more efficient with the proactive assistant, many expressed a preference for less proactivity. This suggests a delicate balance between offering assistance and allowing users to maintain control over their workflow. When the proactive assistant became too intrusive, users tended to ignore its input in favor of their own strategies. However, reducing proactivity too much could limit the assistant’s ability to prompt new discoveries and guide users toward exploring more diverse chart types. Our findings suggest that the proactive assistant positively impacted both task efficiency and the depth of user exploration. As technology continues to advance, there is an opportunity to refine natural language systems to more closely emulate human communication, as described by linguistic theories like Pragmatics. This could enable more intuitive and seamless interactions between users and digital systems.
8. Acknowledgments
This work was supported in part by National Science Foundation awards 2149133 and 2004014.